Field
Report
April
13, 2003
Salar Grande, Atacama Desert, Chile
Agenda
- Autonomous navigation and power measurement experiments
Status and Progress
- Drove Hyperion. We continued basic navigation experiments with Hyperion, ranging farther afield. Today only the
navigation cameras were active, meaning the firewire bus and camera server were disabled. This system worked well
except for occasional sun flare on the lenses late in the day. Problems with the 8 camera configuration were not
resolved although it appears to be a software rather than hardware problem. A new feature was enabled today to
periodically transmit a navigation camera image over the operator interface to monitor performance of cameras.
- Stalled Hyperion. Hyperion climbed in and out the fault valley today several times today and actually blew a
fuse--Hyperion's amplifiers current limit but fuses protect them and it now appears the fuses are too conservative.
The slope was 12°, soft sand. The fuse was replaced and Hyperion continued up the slope.
- Installed panoramic camera. The panoramic camera was installed on the rear mast of the robot. This panoramic camera
as well as a side-looking camera will be used to collect data to develop algorithms for visual odometry.
- Wired fluorescence imager. The underbelly fluorescence imager was installed during vehicle assembly but now the
power wiring has been completed for this prototype instrument. A bit more work is required to control the
notch-filtered LEDs that will be used to illuminate the terrain (so that notch-filtered cameras can look for
fluorescence in a particular wavelength.) This prototype will be used to test strategies for seeking photosynthetic
organisms on the surface.
- Researchers arrive. The second team arrived today with no difficulty. They got set up in base camp and quickly
became involved in various aspects of setting up rover hardware and monitoring operations.
- Modifying solar experiment. We continue to improve the solar experiment, first by increasing the sampling density to
better capture the nonlinear performance curve and next to reduce the amount of oversampling in the linear portions.
It currently takes about 7 minutes to collect all the data for a given direction.
- Voice and data communication working. The voice communication (VHF radios) system is working well with the 10W base
station and high gain antenna and hand-held batteries fully charged. Likewise the rover data communication link is
covering the operational test area.
- Dust devil encountered. Yesterday late afternoon we had a dust devil meander through camp. The vortex was about 50cm
in diameter and it was clearly lifting dust 20-40 meters up into the air. We scrambled for cameras but it ran into the
Chariot del Fuego and dissipated.
- Butterfly encountered. This afternoon a Monarch (or similar) butterfly came flitting through camp. It caused less
dust disturbance than the dust devil. Here, anyway.
Upcoming
- Autonomous navigation and power measurement experiments
- Terrain data collection
- Side-view image data collection
- SPI calibration
- Verify the site DEM and mission planner software updates
- Collect sun sensor test image
Weather
- Night: Clear, Cold
- Day: Clear (no early fog), Windy
This graph shows relative humidity versus hour beginning midnight April 10. The fog event that occurred beginning
midnight April 11 is clearly visible at hour 24 as humidity rises to near 100% yet by midday the humidity is down to a
low of 20%. We'll continue processing this weather data (and improving our graphs) for future reports.
13-Apr-03.humidity.pdf
13-Apr-03.insolation.pdf
13-Apr-03.temperature.pdf
Vulture Count: 15
Quote of the Day
"Wow, trucks full of sulfuric acid, isn't that dangerous?"